OpenStudy (asapbleh):

3(1+x)^1/3 - x(1+x)^-2/3 / (1+x)^2/3 I need help. Please show steps please?

OpenStudy (asapbleh):

\[3(1+x)^{2/3}-x(x+1)^{-2/3}/(1+x)^{2/3}\]

OpenStudy (asapbleh):

they cant show the negative in x(x+1)^-2/3. can anyone please help me solve it?

OpenStudy (mertsj):

\[3(1+x)^{\frac{2}{3}}-\frac{x(x+1)^{-\frac{2}{3}}}{(1+x)^{\frac{2}{3}}}\]

OpenStudy (anonymous):

\[3(1+x)^{\frac{ 2 }{ 3 }}-\frac{ x(x+1)^{-\frac{ 2 }{ 3 }} }{ (1+x)^{\frac{ 2 }{ 3 }} }\] =\[\frac{ 3(1+x)^{\frac{ 2 }{ 3 }+\frac{ 2 }{ 3 }}-x(x+1)^{-\frac{ 2 }{ 3 }} }{ (1+x)^{\frac{ 2 }{ 3 }} }\]

OpenStudy (mertsj):

That is correct so far.

OpenStudy (asapbleh):

sorry if you misunderstood, but the whole thing is divided by (1+x)^2/3. sorry if that was not clear.

OpenStudy (anonymous):

AH okay. @Mertsj You do that for @asapbleh

OpenStudy (anonymous):

I'm going to take a rest.

OpenStudy (mertsj):

\[3(1+x)^{-\frac{1}{3}}-x(1+x)^{-\frac{4}{3}}=(1+x)^{-\frac{4}{3}}[3(1+x)^{\frac{3}{3}}-x]\]

OpenStudy (mertsj):

\[(1+x)^{-\frac{4}{3}}(3+3x-x)=(1+x)^{-\frac{4}{3}}(3+2x)\]

OpenStudy (mertsj):

What were you supposed to do with this?

OpenStudy (asapbleh):

It was a homework problem. it says "simplify the expression. (This type of expression arises in calculus when using the quotient rule."

OpenStudy (mertsj):

There are several different forms you could put it in. That one is probably as good as any.

OpenStudy (asapbleh):

ok, thankyou very much!

OpenStudy (mertsj):

yw